# VidCRAFT3: Camera, Object, and Lighting Control for Image-to-Video Generation

### v3

@@ 4.1 Experimental Setup
Implementation Details. Our model builds upon DynamiCrafter, initialized with its pre-trained weights. During training and inference, video clips are processed at $320\times 512$ resolution with 25 frames. To optimize the model, we employ the Adam optimizer with a learning rate of $1\times 10^{-5}$ and a batch size of 96. The training is conducted on 8 NVIDIA H100 GPUs. For inference, we adopt the DDIM sampler and classifier-free guidance to enhance generation quality and multimodal consistency. During inference, the model averagely uses 20 GB of GPU memory and takes 42 seconds.
@@
CameraCtrl 97.99 96.11 29.41 4.19
MotionCtrl 103.82 188.93 30.18 4.23
CamI2V 98.54 85.03 30.37 4.24
@@
Table 1: Quantitative comparison with SOTA methods on RealEstate10K.
Evaluation Datasets. We evaluate our model on three domain-specific datasets and a generalized test set. For camera motion control, we sample 1,000 samples from the RealEstate10K test set. For object motion control, we sample 1,000 samples from the WebVid-10M test set. For lighting direction control, we sample 1,000 samples from the VLD dataset, covering a wide range of lighting directions. To provide a broader evaluation, we create a generalized test set consisting of 100 videos sourced from copyright-free websites like Pixabay and Pexels, as well as videos generated by T2V models. This dataset spans categories such as human activities, animals, vehicles, indoor scenes, artworks, natural landscapes, and AI-generated images.
4.2 Comparisons with State-of-the-Art Methods
Due to the lack of open-source I2V methods capable of simultaneously controlling camera and object motion, we evaluate VidCRAFT3 separately for camera motion control and object motion control, comparing it against SOTA methods in each domain.
@@ 4.2 Comparisons with State-of-the-Art Methods
Image Conductor 150.26 242.01 29.69 12.96
Motion-I2V 128.35 171.35 30.92 3.96
@@
Table 2: Quantitative comparison with SOTA methods on WebVid-10M.
Camera Motion Control. VidCRAFT3 demonstrates superior performance in camera motion control compared to state-of-the-art (SOTA) methods like CameraCtrl, CamI2V, and MotionCtrl on the RealEstate 10K dataset. As shown in Table 1 and Fig. 5, VidCRAFT3 excels in camera motion control, as evidenced by its CamMC score of 4.07, which is lower than CameraCtrl (4.19), CamI2V (4.24), and MotionCtrl (4.23). VidCRAFT3 also achieves better results in other key metrics indicate that VidCRAFT3 not only controls camera motion more accurately but also generates videos with higher visual quality, temporal coherence, and semantic alignment. Qualitatively, VidCRAFT3 generates smoother and more realistic camera movements with fewer artifacts, particularly in complex scenes. This success stems from the Image2Cloud module, which enable precise 3D scene reconstruction and fine-grained camera control.
Object Motion Control. VidCRAFT3 demonstrates exceptional performance in object motion control, as highlighted in Table 2. The model achieves an ObjMC score of 3.51, which is lower than Image Conductor (12.96) and Motion-I2V (3.96). This indicates that VidCRAFT3 more accurately aligns the generated object trajectories with the ground truth, resulting in more realistic and faithful object motion. Additionally, VidCRAFT3 outperforms in other key metrics, showcasing superior visual quality, temporal coherence, and semantic alignment. Qualitatively, as shown in Fig. 6, VidCRAFT3 generates more realistic and consistent object movements compared to Image Conductor and Motion-I2V. The model effectively captures the dynamics of object motion, ensuring smooth transitions and natural interactions within the scene. These results highlight VidCRAFT3’s robust object motion control, driven by its advanced ObjMotionNet, which effectively captures and controls complex motion patterns.
@@
CameraCtrl 3.4% 5.4% 4.7%
MotionCtrl 5.7% 7.3% 6.5%
CamI2V 5.6% 6.8% 5.2%
@@
Image Conductor 10.3% 11.9% 11.2%
Motion-I2V 14.8% 10.9% 13.2%
@@
Table 3: User study. Results demonstrating our method’s superior performance in camera and object motion control compared to baseline approaches across all metrics.
@@
Text Cross-Atten 111.08 121.95 22.77 18.14 0.72 0.13 5.31
Time Embed. 101.71 123.31 22.70 19.07 0.73 0.12 5.21
Lighting Cross-Attn 100.83 117.69 23.70 19.49 0.74 0.11 5.00
Table 4: Ablation of lighting embedding integration strategies on VLD. Text Cross-Atten method concatenates lighting embedding with text embedding and integrates them into the model through text cross-attention. Time Embed method adds lighting embedding to time embedding. The proposed Lighting Cross-Atten method, detailed in Sec. 3.2, demonstrates enhanced performance in integrating lighting conditions.
User Study. To evaluate our method, we conducted a user study with 10 participants assessing 200 randomly selected results from our method and five baselines. The participants evaluated the results based on four key metrics: Camera Motion Precision, Object Motion Precision, Visual Quality, and Overall Quality. As shown in Table 3, our method achieved over 80% in camera motion control and over 74% in object motion control across all metrics, demonstrating precise and visually appealing control, validating its effectiveness and robustness in real-world applications.
4.3 Ablation Study
Training Strategy. We compare training with Dense, Sparse, and Dense+Sparse trajectories. As shown in Table 5, Dense training (ObjMC: 4.39) provides rich motion information but struggles during inference due to sparse inputs. Sparse training (ObjMC: 4.05) improves inference alignment but lacks motion detail. The Dense+Sparse approach achieves the best results (ObjMC: 3.51), leveraging dense trajectories for robust learning and fine-tuning with sparse trajectories for better generalization. This hybrid strategy enhances object motion control and video quality, as shown by superior FID (87.12) and FVD (120.65) scores.
Lighting Embedding Integration Strategies. We compare different lighting embedding integration strategies. As shown in Table 4, Lighting Cross-Atten achieves the best results, with an PSNR of 19.49, SSIM of 0.74, and LPIPS of 0.11, outperforming other methods. These findings confirm the effectiveness of Lighting Cross-Attention for precise lighting direction control.
Representation of Lighting Direction. We compare Fourier Embedding [31] and SH Encoding for representation of lighting direction. As shown in Table 6, SH Encoding outperforms Fourier Embedding, achieving better PSNR (19.49), SSIM (0.74), and LPIPS (0.11). These results confirm SH Encoding’s effectiveness for precise lighting control in image-to-video generation. For more results, please refer to the supplementary materials.
@@ 4.3 Ablation Study
Dense+Sparse 87.12 120.65 32.99 3.51
Table 5: Ablation of training strategy on WebVid-10M.
@@
Table 6: Ablation of representation of lighting direction on VLD.
5 Conclusions
We present VidCRAFT3, a novel framework for precise image-to-video generation with simultaneous control over camera motion, object motion, and lighting direction. It integrates Image2Cloud for 3D point cloud generation, ObjMotionNet for encoding sparse object trajectories, and a Spatial Triple-Attention Transformer for incorporating lighting direction. The VideoLightingDirection dataset and a three-stage training strategy address real-world annotation challenges. Experiments show that VidCRAFT3 outperforms state-of-the-art methods in control granularity and visual coherence, advancing video generation toward more realistic and versatile applications.
Supplementary Material
Appendix A Preliminary
Image-to-video generation. Given a reference image $I_{\text{ref}}$ and some text description, image-to-video generation aims to generate a sequence of video frames $I=\{I_{1},I_{2},\ldots,I_{F}\}$ with $I_{1}=I$. In this work, we also allow the user to provide three forms of control signals: camera motion, object motion, and lighting direction. The video frames will be generated according to the text description and all the control signals provided. Specifically, the camera motion is represented by the movement of a set of camera extrinsics $E=\{E_{1},E_{2},\ldots,E_{F}\}$; the object motion is represented by a few sparse sequences of image pixels $\{s_{1}^{1},s_{1}^{2},s_{1}^{3},\ldots,s_{1}^{F}\},\{s_{1}^{2},s_{2}^{2},s_{2}^{3},\ldots,s_{2}^{F}\},\ldots$; the lighting direction is simply a unit 3d vector $L\in\boldsymbol{R}^{3}$.
Video Diffusion model (VDMs) represent a class of generative models that extend the principles of image diffusion to the domain of video generation. These models operate by defining a forward diffusion process that gradually transforms an initial video sample $x_{0}\sim p_{\text{data}}(x)$ into Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$ over $T$ timesteps. The reverse process, parameterized by a denoising network $\epsilon_{\theta}(x_{t},t)$, learns to iteratively denoise the noisy latent representation $x_{t}$ to recover the original data $x_{0}$. The training objective is formulated as:
$\min_{\theta}\mathbb{E}_{t,x\sim p_{\text{data}},\epsilon\sim\mathcal{N}(0,I)}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2},$
To address the computational challenges associated with high-dimensional video data, Latent Diffusion Models (LDMs) are often employed. In this framework, a video $x\in\mathbb{R}^{F\times 3\times H\times W}$ is first encoded into a lower-dimensional latent space $\mathbf{z}=\mathcal{E}(x)$, where $\mathbf{z}\in\mathbb{R}^{F\times C\times h\times w}$. The diffusion and denoising processes are then performed in this latent space, significantly reducing computational complexity. The denoising process is conditioned on additional inputs $\mathbf{c}$, such as text prompts or motion control signals, enabling the generation of videos that adhere to specific semantic or temporal constraints. The final video is reconstructed through a decoder $\hat{x}=\mathcal{D}(\mathbf{z})$, ensuring both spatial and temporal coherence in the generated output.
Appendix B VLD Dataset Samples
Fig. 7 illustrates representative samples from our proposed VideoLightingDirection (VLD) Dataset, a synthetic dataset specifically designed to model complex light-object interactions. The dataset comprises two types of scenes: (a) Haven scenes, in which 3D models from Poly Haven are placed centrally within HDR environments; and (b) BOP scenes, featuring randomly positioned BOP models in six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions while maintaining a consistent camera trajectory. This design highlights how changes in lighting direction impact object shading, reflections, and overall visual coherence, providing valuable data for training and evaluating models requiring precise lighting direction annotations.
Figure 7: Illustrations of samples from the proposed VideoLightingDirection (VLD) Dataset, featuring synthetic scenes designed to model complex light-object interactions. (a) Haven scenes, where 3D models from Poly Haven are placed at the center of HDR environments. (b) BOP scenes, with BOP models randomly positioned within six-plane textured rooms to simulate indirect lighting. Each subset includes video frames captured under two distinct lighting conditions, maintaining consistent camera trajectories.
Appendix C Qualitative Results of Ablation Study
@@
Training Strategy. Fig. 8 qualitatively compares videos generated using Dense, Sparse, and Dense+Sparse training strategies. Dense training results exhibit detailed object motion but struggle with precise alignment when tested with sparse trajectories. Sparse training achieves better alignment but at the cost of reduced motion detail. The Dense+Sparse approach effectively combines the strengths of both, delivering superior motion coherence and overall visual quality.
Lighting Embedding Integration Strategies. Fig. 9 compares the generated videos of different lighting embedding integration strategies. Lighting Cross-Attention demonstrates clear superiority, providing precise control over lighting directions, resulting in more realistic shadows, reflections, and overall lighting effects compared to Text Cross-Attention and Time Embedding.
Representation of Lighting Direction. Fig. 10 presents a visual comparison between Fourier Embedding and SH Encoding for representing lighting direction. SH Encoding consistently produces more realistic and detailed lighting effects, including enhanced shading and reflections, further supporting the quantitative findings in the main text.
@@
Appendix D Additional Qualitative Results
This section provides additional qualitative comparisons illustrating VidCRAFT3’s capabilities across multiple control dimensions.
Camera Motion + Lighting Direction Control. Fig. 11 demonstrates VidCRAFT3’s effectiveness in simultaneously controlling camera motion and lighting direction, showing clear advantages in lighting realism and trajectory accuracy compared to original sequences.
Camera Motion + Object Motion Control. Fig. 12 presents examples of simultaneous control over camera and object motions. VidCRAFT3 accurately reproduces intended object trajectories and maintains specified camera movements, closely matching the ground truth.
Camera Motion Control. Fig. 13 compares VidCRAFT3’s results with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth sequences, highlighting superior camera trajectory accuracy and visual coherence.
Object Motion Control. Fig. 14 provides comparisons against Image Conductor and Motion-I2V, showcasing VidCRAFT3’s superior capability in generating realistic object trajectories and visually coherent motion across diverse scenarios.
Lighting Direction Control. Fig. 15 further illustrates VidCRAFT3’s precise control over lighting direction, achieving accurate shading, shadows, and reflections, consistent with ground-truth sequences.
@@ Appendix D Additional Qualitative Results
Appendix E Limitations
VidCRAFT3 can generate unstable results for many specific situations, e.g., large human motion, phyisical interactions, significant change of lighting conditions. This is mostly due to the lack of diverse training data and we believe the neural architecture can also be improved to promote a better understanding of physics and 3D spatial relationships of objects. The camera motion control and object motion control datasets contain inaccurate annotations of camera pose and motion trajectories, which sometimes result in blurred generated videos. Currently, VidCRAFT3 only offers control over the lighting direction and therefore can be improved to support full HDR or other more fine-grained representations of the light field.
Figure 8: Qualitative comparisons of training strategies (Dense, Sparse, Dense+Sparse) in the ablation study conducted on WebVid-10M. Results clearly illustrate that Dense training struggles with alignment given sparse input trajectories during inference, Sparse training improves alignment but loses motion detail, while Dense+Sparse training effectively combines both strategies, resulting in superior alignment and enhanced motion realism compared to the ground truth (GT).
Figure 9: Qualitative comparison of lighting embedding integration strategies (Text Cross-Attn, Time Embedding, Lighting Cross-Attn) in the ablation study on the VLD dataset. Results clearly demonstrate that the Lighting Cross-Attn strategy provides more accurate and realistic lighting control, achieving superior shadows, reflections, and overall visual fidelity compared to Text Cross-Attn and Time Embedding, aligning closely with the ground truth (GT).
Figure 10: Qualitative comparison between Fourier Embedding and SH Encoding for representing lighting direction on the VLD dataset. Results demonstrate that SH Encoding consistently produces more realistic lighting effects, achieving better shading accuracy and reflection details compared to Fourier Embedding, aligning closely with the ground truth (GT).
Figure 11: Additional qualitative results demonstrating simultaneous control over camera motion and lighting direction. Each pair of rows compares the original video with the corresponding relighting results generated by VidCRAFT3, highlighting the framework’s effectiveness in precisely adjusting lighting direction while preserving accurate camera motion.
Figure 12: Additional qualitative results showcasing simultaneous camera motion and object motion control. Each example compares our generated video (Ours) with ground-truth videos (GT), demonstrating VidCRAFT3’s capability to accurately reproduce intended object motion while precisely maintaining the specified camera motion.
@@ Appendix E Limitations
Figure 15: Additional qualitative results demonstrating lighting direction control. Our generated videos (”Ours”) are compared against ground-truth sequences (”GT”), clearly highlighting VidCRAFT3’s capability to accurately reproduce different lighting direction, achieving realistic shading, shadowing, and reflection details across diverse scenes.

### v4

@@ IV-A Experimental Setup
Implementation Details. Our model builds upon DynamiCrafter, initialized with its pre-trained weights. During training and inference, video clips are processed at $320\times 512$ resolution with 25 frames. We optimize the model using Adam with a learning rate of $1\times 10^{-5}$ and a global batch size of 96, and adopt classifier-free training with an unconditional drop probability of $0.05$. Training is conducted on 8 NVIDIA H100 GPUs. For inference, we use a DDIM sampler and classifier-free guidance with a guidance scale of $7.5$. On average, inference uses about 20 GB of GPU memory and takes roughly 42 seconds per sample.
Evaluation Datasets. We evaluate our model on three domain-specific datasets and a generalized test set. For camera motion control, we sample 1,000 samples from the RealEstate10K test set. For object motion control, we sample 1,000 samples from the WebVid-10M test set. For lighting direction control, we evaluate on 1,000 samples from VLD, stratified as 500 Haven-type scenes and 500 BOP-type scenes, covering a wide range of lighting directions. To provide a broader evaluation, we create a generalized test set consisting of 100 videos sourced from copyright-free websites such as Pixabay and Pexels, as well as videos generated by T2V models. This dataset spans categories such as human activities, animals, vehicles, indoor scenes, artworks, natural landscapes, and AI-generated images.
Figure 9: Qualitative results of the same reference image under different camera motion, object motion and lighting direction. For each subfigure, we vary only one control while keeping the other two fixed: (a) camera motion, (b) object motion, and (c) lighting direction. The leftmost panels visualize the target control (camera trajectory / object trajectory / lighting direction). VidCRAFT3 faithfully follows the control signal while preserving content fidelity.”
IV-B Main Results
VidCRAFT3 demonstrates strong controllability and visual fidelity across camera motion, object motion, and lighting direction, while preserving temporal and semantic consistency. As no open-source I2V method supports simultaneous camera–object control, we evaluate VidCRAFT3 separately on camera motion and object motion against state-of-the-art baselines, then present qualitative joint control results.
TABLE I: Quantitative comparison of camera motion control on RealEstate10K. VidCRAFT3 shows stronger camera adherence with improved visual fidelity, temporal coherence, and semantic alignment.
@@ IV-B Main Results
CameraCtrl [7] 97.99 96.11 29.41 4.19
MotionCtrl [6] 103.82 188.93 30.18 4.23
CamI2V [33] 98.54 85.03 30.37 4.24
@@
TABLE II: Quantitative comparison of object motion control on WebVid-10M. VidCRAFT3 shows stronger object adherence and identity preservation, with improved visual fidelity, temporal coherence, and semantic alignment
@@
Image Conductor [55] 150.26 242.01 29.69 12.96
Motion-I2V [58] 128.35 171.35 30.92 3.96
@@
IV-B1 Camera Motion Control
VidCRAFT3 achieves accurate and stable camera motion control on RealEstate10K. As shown in Fig. 9(a), under different complex camera trajectory controls, the model accurately generates videos with the specified camera motion while preserving scene content. Quantitatively, Table I reports the best results across all metrics: CamMC 4.07 (better than the best baseline, 4.19), FID 75.62, FVD 49.77, and CLIPSIM 32.32, outperforming CameraCtrl, CamI2V, and MotionCtrl. These improvements indicate more precise camera control, along with better visual quality, temporal coherence, and semantic alignment. Qualitatively, Fig. 7 shows smoother and more realistic camera motion with fewer artifacts, particularly in complex scenes. We attribute these improvements to the explicit 3D priors provided by Image2Cloud, which anchor camera motion to a consistent scene geometry and significantly reduce drift during large viewpoint changes.
IV-B2 Object Motion Control
VidCRAFT3 demonstrates exceptional performance in object motion control on WebVid-10M. As shown in Fig. 9(b), under different complex object trajectories controls, the model can accurately control the object motion in the video. Quantitatively, Table II reports that the model achieves an ObjMC score of 3.51, which is lower than Image Conductor (12.96) and Motion-I2V (3.96). This indicates that VidCRAFT3 achieves closer alignment between generated and target trajectories, resulting in more realistic and faithful object motion. Additionally, VidCRAFT3 outperforms other methods in other metrics, showcasing superior visual quality, temporal coherence, and semantic alignment. Qualitatively, as shown in Fig. 8, VidCRAFT3 generates more realistic and consistent object movements compared to Image Conductor and Motion-I2V. The model effectively captures the dynamics of object motion, ensuring smooth transitions and natural interactions within the scene. These results highlight VidCRAFT3’s robust object motion control, driven by its advanced ObjMotionNet, which effectively captures and controls complex motion patterns.
TABLE III: Quantitative results for lighting direction control on VLD across 1,000 clips, stratified by scene type: 500 Haven-type and 500 BOP-type.
@@ IV-B2 Object Motion Control
Haven-type 105.70 123.65 20.45 17.79 0.67 0.13 5.21
BOP-type 95.96 111.73 26.95 21.19 0.81 0.09 4.79
Overall 100.83 117.69 23.70 19.49 0.74 0.11 5.00
TABLE IV: User study. Results demonstrating our method’s superior performance in camera and object motion control compared to baseline approaches across all metrics.
[…]
CameraCtrl [7] 3.4% 5.4% 4.7%
MotionCtrl [6] 5.7% 7.3% 6.5%
CamI2V [33] 5.6% 6.8% 5.2%
@@
Image Conductor [55] 10.3% 11.9% 11.2%
Motion-I2V [58] 14.8% 10.9% 13.2%
@@
Figure 10: Qualitative results demonstrating simultaneous control over camera motion and object motion. For each block, we show the specified camera trajectory and object trajectories as controls, followed by frames from the ground-truth video (GT, top) and from VidCRAFT3 (Ours, bottom).
IV-B3 Lighting Direction Control
VidCRAFT3 enables accurate and realistic lighting direction control for both synthetic and real reference images. As shown in Fig. 9(c), with camera and object motion fixed, supplying a sequence of lighting directions produces videos whose illumination follows the control signal while preserving appearance. Notably, although training samples contain only static lighting directions, the model produces coherent responses to time-varying direction sequences at test time, demonstrating the combined generalization of direction-level conditioning and temporal modeling. Prior methods largely target portraits and rely on HDR environment maps rather than interactive, direction-level control in general scenes. Their evaluation protocols are also incompatible with our VLD benchmark. Accordingly, we report results on VLD and verify design choices via ablations (Table VIII, Table VII). Table III presents metrics on the Haven and BOP scene types: VidCRAFT3 performs comparably on both, with slightly better results on BOP. This suggests that a single spotlight is easier to disentangle than environment light + spotlight, while the small gap indicates that our model exhibits scene-agnostic domain generalization.
IV-B4 Joint Control
VidCRAFT3 delivers accurate and coherent joint control of camera motion, object motion, and lighting direction. We use the same inference setup as in the single-control setting and compose a subset of controls in a single forward pass. Qualitative results are shown in Fig. 1, Fig. 12, Fig. 11, and Fig. 10. For camera + lighting (Fig. 12, Fig. 11), the model follows the prescribed camera trajectory while maintaining consistent shading, highlights, and cast shadows across viewpoints, indicating temporally stable lighting. For camera + object (Fig. 10), the model preserves the global camera trajectory and adheres to sparse object trajectories, retaining identity and shape with parallax that matches the camera motion. With all three controls (Fig. 1), object motion aligns with target trajectories, lighting changes remain coherent with viewpoint transitions, and the scene layout stays stable under large viewpoint changes. These results demonstrate clean compositionality without interference among controls and sustained temporal consistency across frames.
Figure 11: Qualitative results demonstrating simultaneous control over camera motion and lighting direction. For each row, we show a reference image, the specified camera trajectory and lighting direction controls, and the relighting video frames generated by VidCRAFT3.
Figure 12: Qualitative results demonstrating simultaneous control over camera motion and lighting direction on VLD. The top block shows a Haven-type scene and the bottom block shows a BOP-type scene. Our generated videos (”Ours”) are compared against ground-truth sequences (”GT”).
IV-C User Study
To evaluate VidCRAFT3, we conducted a user study with 10 participants assessing 200 randomly selected results, comprising 100 camera motion samples (50 from copyright-free and 50 from the RealEstate10K test split) and 100 object motion samples (50 from copyright-free and 50 from the WebVid-10M test split), drawn from the outputs of our method and five baselines. The participants evaluated the results based on four metrics: Camera Motion Precision, Object Motion Precision, Visual Quality, and Overall Quality. As shown in Table IV, our method achieved over 80% in camera motion control and over 74% in object motion control across all metrics, demonstrating precise and visually appealing control, validating its effectiveness and robustness in real-world applications.
IV-D Ablation Studies
@@
[…]
Full three-Stage (S1 $\rightarrow$ S2 $\rightarrow$ S3)
Stage 1 (camera) 3.98
Stage 2 (object-dense + VLD) 4.19 4.39 18.21 0.60 0.20
Stage 3 (object-sparse + VLD) 4.07 3.51 19.49 0.74 0.11
No-S1 (S2 $\rightarrow$ S3)
Stage 2 (dense + VLD) 6.52 4.35 18.47 0.68 0.21
Stage 3 (sparse + VLD) 6.17 3.45 19.34 0.81 0.10
[…]
@@
[…]
Training Strategy. To validate the effectiveness of our three-stage training strategy, we conduct two experiments: (1) whether the progressive three-stage training strategy affects the capabilities learned in the previous stages (e.g., camera motion control), and (2) whether the camera prior learned in Stage 1 is necessary. As shown in Table V, in the full three-stage advancing from Stage 2 to Stage 3 improves object adherence and relighting quality while also improving camera adherence. In contrast, the No-S1 variant achieves similar object and relighting scores at Stage 3 but shows a pronounced drop in CamMC, indicating that Stage 1 provides a crucial camera and geometry prior.
[…]
Lighting Embedding Integration Strategies. We compare different lighting embedding integration strategies. (1) Text Cross-Attn method concatenates lighting embedding with text embedding and integrates them into the model through text cross-attention. (2) Time Embed method adds lighting embedding to time embedding. (3) The proposed Lighting Cross-Atten method, which introduces a dedicated lighting cross-attention branch inside the Spatial Triple-Attention Transformer. As shown in Table VII, Lighting Cross-Attn achieves the best results and outperforms the other methods across all metrics, indicating that explicit, decoupled lighting attention integrates lighting directions more effectively than routing them through text or time embedding.
Representation of Lighting Direction. We compare Fourier Embedding [104] and SH Encoding for representation of lighting direction. The resulting lighting embeddings are fed to the Lighting Cross-Attn branch. As shown in Table VIII, SH Encoding outperforms Fourier Embedding across all metrics on VLD. We attribute these gains to the smooth, rotation-aware angular basis of SH, which provides a more stable and geometry-consistent signal for lighting than Fourier features. For more results, please refer to the supplementary materials.
TABLE VII: Ablation of lighting embedding integration strategies on VLD. The proposed Lighting Cross-Attn achieves the best overall quality and lighting control.
[…]
Text Cross-Attn 111.08 121.95 22.77 18.14 0.72 0.13 5.31
Time Embed. 101.71 123.31 22.70 19.07 0.73 0.12 5.21
Lighting Cross-Attn 100.83 117.69 23.70 19.49 0.74 0.11 5.00
TABLE VIII: Ablation of representation of lighting direction on VLD. SH Encoding surpasses Fourier Embedding across fidelity and alignment metrics.
@@
V Conclusions
In conclusion, we present VidCRAFT3, a unified and flexible framework for precise image-to-video generation that enables simultaneous control over camera motion, object motion, and lighting direction. VidCRAFT3 integrates three core components: Image2Cloud, which generates 3D point clouds from reference images; ObjMotionNet, which encodes sparse object trajectories using multi-scale optical flow features; and Spatial Triple-Attention Transformer, which incorporates lighting direction embeddings through parallel cross-attention modules. The introduction of the VideoLightingDirection dataset, combined with a three-stage training strategy, effectively addresses the challenges posed by the lack of annotated real-world datasets. Extensive experiments demonstrate that VidCRAFT3 produces high-quality video content, outperforming state-of-the-art methods in terms of control granularity and visual coherence. Our work advances the field of video generation by enabling enhanced control over multiple visual elements, thereby paving the way for more realistic and versatile applications.
Acknowledgments
This paper is supported by the Doubao Fund.
Qualitative Results of Ablation Study
@@ Acknowledgments
[…]
Lighting Embedding Integration Strategies. Fig. 14 qualitatively compares three lighting embedding integration strategies: Text Cross-Attention, Time Embedding, and Lighting Cross-Attention. Text Cross-Attention integrates lighting embeddings with text embeddings through cross-attention, resulting in less accurate lighting control and suboptimal lighting effects. Time Embedding improves upon this by incorporating lighting information into the time embedding, providing better consistency but still lacking in precision. Lighting Cross-Attention excels by directly integrating lighting embeddings through dedicated cross-attention, offering superior control over lighting directions and producing more realistic shadows, reflections, and overall lighting effects. This strategy aligns more closely with the ground truth (GT) and achieves better visual fidelity (cf. Table VII).
Representation of Lighting Direction. Fig. 15 qualitatively compares two methods for representing lighting direction: Fourier Embedding and SH Encoding. Fourier Embedding represents lighting direction using periodic basis functions but struggles with accurately capturing complex lighting effects. In contrast, SH Encoding leverages spherical harmonics to better model the angular properties of lighting, producing more realistic and detailed lighting effects, such as enhanced shading and reflections. These improvements result in outputs that align more closely with the ground truth (GT), as demonstrated in both visual and quantitative comparisons (cf. Table VIII).
@@
Additional Qualitative Results
This section presents additional qualitative comparisons, highlighting VidCRAFT3’s capabilities in controlling camera motion and object motion.
Camera Motion Control. Fig. 16 compares VidCRAFT3’s results with state-of-the-art methods (CameraCtrl, MotionCtrl, CamI2V) and ground-truth sequences, highlighting superior camera trajectory accuracy and visual coherence. These examples demonstrate VidCRAFT3’s superior capability in generating visually coherent and precise camera motion across diverse scenarios.
Object Motion Control. Fig. 17 provides comparisons against Image Conductor and Motion-I2V, showcasing VidCRAFT3’s superior capability in generating realistic object trajectories and visually coherent motion across diverse scenarios. The examples demonstrate VidCRAFT3’s improved capability in accurately reproducing specified object trajectories, achieving more realistic and coherent object motion across diverse scenarios.
@@
Limitations
VidCRAFT3 can generate unstable results for many specific scenarios, e.g., large human motion, physical interactions, and significant changes of lighting conditions. This is mainly due to the lack of diverse training data. Additionally, we believe the neural architecture can be further improved to enhance the model’s understanding of physics and 3D spatial relationships. Inaccurate annotations in the camera motion control and object motion control datasets, particularly for camera poses and motion trajectories, can sometimes lead to blurred videos. Currently, VidCRAFT3 only offers control over lighting direction. To improve, the model could be extended to support full HDR or other more fine-grained representations of the light field.
[…]
Figure 14: Qualitative comparison of lighting embedding integration strategies (Text Cross-Attn, Time Embedding, and Lighting Cross-Attn) in the ablation study on the VLD dataset. Each example shows frames generated under different lighting embedding strategies. Text Cross-Attn integrates lighting embeddings with text embeddings through cross-attention but results in less accurate lighting control. Time Embedding improves consistency by incorporating lighting information into time embedding, but still lacks precision. Lighting Cross-Attn directly integrates lighting embeddings with dedicated cross-attention, providing superior control over lighting directions, producing more realistic shadows, reflections, and overall lighting effects, and aligning more closely with the ground truth (GT) compared to the other strategies.
Figure 15: Qualitative comparison between Fourier Embedding and SH Encoding for representing lighting direction in the ablation study on the VLD dataset. Each example shows frames generated under different lighting direction representations. Fourier Embedding represents lighting direction using periodic basis functions but struggles with accurately capturing complex lighting effects. SH Encoding utilizes spherical harmonics to better model lighting direction, producing more realistic and detailed lighting effects, including enhanced shading and reflections. SH Encoding aligns more closely with the ground truth (GT), providing superior visual fidelity and more accurate lighting control compared to Fourier Embedding.